OpenStudy (anonymous):

The difference between -512 and 431 is __________ . -98 -843 121 -943

OpenStudy (anonymous):

-943

OpenStudy (johnn.):

-512−431 =−512−431 =−512+−431 =−943

OpenStudy (anonymous):

Thank you both @imdestroyer69 @johndp1236

OpenStudy (anonymous):

np :)